# A Comprehensive Survey on World Models for Embodied AI

### v1

@@ 5.2 Scene Understanding
OccWorld-O[93] Occ None Pred. 66.38 25.78 15.14 10.51 17.14 62.29 34.63 25.07 20.18 26.63
OccLLaMA-O[18] Occ None Pred. 75.20 25.05 19.49 15.26 19.93 63.76 34.56 28.53 24.41 29.17
RenderWorld-O[156] Occ None Pred. - 28.69 18.89 14.83 20.80 - 37.74 28.41 24.08 30.08
DTT-O[98] Occ None Pred. 85.50 37.69 29.77 25.10 30.85 92.07 76.60 74.44 72.71 74.58
DFIT-OccWorld-O[174] Occ None Pred. - 31.68 21.29 15.18 22.71 - 40.28 31.24 25.29 32.27
COME-O[213] Occ None Pred. - 30.57 19.91 13.38 21.29 - 36.96 28.26 21.86 29.03
DOME-O[94] Occ None GT 83.08 35.11 25.89 20.29 27.10 77.25 43.99 35.36 29.74 36.36
COME-O[213] Occ None GT - 42.75 32.97 26.98 34.23 - 50.57 43.47 38.36 44.13
OccWorld-T[93] Camera Semantic LiDAR Pred. 7.21 4.68 3.36 2.63 3.56 10.66 9.32 8.23 7.47 8.34
OccWorld-S[93] Camera None Pred. 0.27 0.28 0.26 0.24 0.26 4.32 5.05 5.01 4.95 5.00
RenderWorld-S[156] Camera None Pred. - 2.83 2.55 2.37 2.58 - 14.61 13.61 12.98 13.73
COME-S[213] Camera None Pred. - 25.57 18.35 13.41 19.11 - 45.36 37.06 30.46 37.63
OccWorld-D[93] Camera Occ Pred. 18.63 11.55 8.10 6.22 8.62 22.88 18.90 16.26 14.43 16.53
OccWorld-F[93] Camera Occ Pred. 20.09 8.03 6.91 3.54 6.16 35.61 23.62 18.13 15.22 18.99
OccLLaMA-F[18] Camera Occ Pred. 37.38 10.34 8.66 6.98 8.66 38.92 25.81 23.19 19.97 22.99
DFIT-OccWorld-F[174] Camera Occ Pred. - 13.38 10.16 7.96 10.50 - 19.18 16.85 15.02 17.02
DTT-F[98] Camera Occ Pred. 43.52 24.87 18.30 15.63 19.60 54.31 38.98 37.45 31.89 36.11
DOME-F[94] Camera Occ GT 75.00 24.12 17.41 13.24 18.25 74.31 35.18 27.90 23.44 28.84
COME-F[213] Camera Occ GT - 26.56 21.73 18.49 22.26 - 48.08 43.84 40.28 44.07
@@ 5.3 Control Tasks
PlaNet[38] 5M 469 496 495 945 333/20
Dreamer[10] 5M 935 895 499 962 823/20
Dreaming[110] 500k 905 566 762 469 610/12
TransDreamer[28] 2M - 865 - 933 893/4
DreamerPro[111] 1M 873 897 811 - 857/6
MWM[41] 1M - 670 - 690/7
HRSSM[25] 500k 910 - 960 - 938/3
DisWM[112] 1M 960 820 - 920 879/5
@@ Methods
Criteria VidMan[55] ManiGaussian[53] ManiGaussian++[80] DreMa[60] TesserAct[78]
@@
UniAD[254] Camera Map & Box & Motion & Tracklets & Occ 0.48 0.96 1.65 1.03 0.05 0.17 0.71 0.31
UniAD+DriveWorld[87] Camera Map & Box & Motion & Tracklets & Occ 0.34 0.67 1.07 0.69 0.04 0.12 0.41 0.19
GenAD[92] Camera Map & Box & Motion 0.36 0.83 1.55 0.91 0.06 0.23 1.00 0.43
FSDrive[101] Camera Map & Box & QA 0.40 0.89 1.60 0.96 0.07 0.12 1.02 0.40
OccWorld-T[93] Camera Semantic LiDAR 0.54 1.36 2.66 1.52 0.12 0.40 1.59 0.70
Doe-1[134] Camera QA 0.50 1.18 2.11 1.26 0.04 0.37 1.19 0.53
SSR[160] Camera None 0.24 0.65 1.36 0.75 0.00 0.10 0.36 0.15
OccWorld-S[93] Camera None 0.67 1.69 3.13 1.83 0.19 1.28 4.59 2.02
Epona[148] Camera None 0.61 1.17 1.98 1.25 0.01 0.22 0.85 0.36
RenderWorld[156] Camera None 0.48 1.30 2.67 1.48 0.14 0.55 2.23 0.97
Drive-OccWorld[157] Camera None 0.32 0.75 1.49 0.85 0.05 0.17 0.64 0.29
OccWorld-D[93] Camera Occ 0.52 1.27 2.41 1.40 0.12 0.40 2.08 0.87
OccWorld-F[93] Camera Occ 0.45 1.33 2.25 1.34 0.08 0.42 1.71 0.73
OccLLaMA-F[18] Camera Occ 0.38 1.07 2.15 1.20 0.06 0.39 1.65 0.70
DTT-F[98] Camera Occ 0.35 1.01 1.89 1.08 0.08 0.33 0.91 0.44
DFIT-OccWorld-V[174] Camera Occ 0.42 1.14 2.19 1.25 0.09 0.19 1.37 0.55
NeMo[161] Camera Occ 0.39 0.74 1.39 0.84 0.00 0.09 0.82 0.30
OccWorld-O[93] Occ None 0.43 1.08 1.99 1.17 0.07 0.38 1.35 0.60
OccLLaMA-O[18] Occ None 0.37 1.02 2.03 1.14 0.04 0.24 1.20 0.49
RenderWorld-O[156] Occ None 0.35 0.91 1.84 1.03 0.05 0.40 1.39 0.61
DTT-O[98] Occ None 0.32 0.91 1.76 1.00 0.08 0.32 0.51 0.30
DFIT-OccWorld-O[174] Occ None 0.38 0.96 1.73 1.02 0.07 0.39 0.90 0.45
  • 1 Note: Results are reported following the UniAD [254] protocol. Method variants are denoted by their input source: O (ground-truth), V (camera-predicted), D (TPVFormer), F (FBOCC), etc.
@@
6 Challenges and Trends
This section reviews the open challenges and emerging directions for world models in embodied AI. We discuss them across three dimensions: §6.1 Data & Evaluation, §6.2 Computational Efficiency, and §6.3 Modeling Strategies.
6.1 Data & Evaluation
@@
Evaluation practices face similar limitations. Metrics such as FID and FVD emphasize pixel fidelity while ignoring physical consistency, dynamics, and causality. Recent benchmarks, such as EWM-Bench [255], introduce new measures but remain task-specific and lack cross-domain standards.
Future Directions.  Recent initiatives such as OpenDV-2K [90] and VideoMix22M [14] highlight the growing focus on large-scale pretraining and broader modality coverage, yet resources remain fragmented and domain specific. Future work should prioritize constructing unified multimodal, cross-domain datasets to enable transferable pretraining, while advancing evaluation frameworks that move beyond perceptual realism to assess physical consistency, causal reasoning, and long-horizon dynamics.
6.2 Computational Efficiency

### v2

@@ V-B Scene Understanding
OccWorld-O[zheng2024occworld] Occ None Pred. 66.38 25.78 15.14 10.51 17.14 62.29 34.63 25.07 20.18 26.63
OccLLaMA-O[wei2024occllama] Occ None Pred. 75.20 25.05 19.49 15.26 19.93 63.76 34.56 28.53 24.41 29.17
RenderWorld-O[yan2025renderworld] Occ None Pred. - 28.69 18.89 14.83 20.80 - 37.74 28.41 24.08 30.08
DTT-O[xu2025delta] Occ None Pred. 85.50 37.69 29.77 25.10 30.85 92.07 76.60 74.44 72.71 74.58
DFIT-OccWorld-O[zhang2024efficient] Occ None Pred. - 31.68 21.29 15.18 22.71 - 40.28 31.24 25.29 32.27
COME-O[shi2025come] Occ None Pred. - 30.57 19.91 13.38 21.29 - 36.96 28.26 21.86 29.03
DOME-O[gu2024dome] Occ None GT 83.08 35.11 25.89 20.29 27.10 77.25 43.99 35.36 29.74 36.36
COME-O[shi2025come] Occ None GT - 42.75 32.97 26.98 34.23 - 50.57 43.47 38.36 44.13
OccWorld-T[zheng2024occworld] Camera Semantic LiDAR Pred. 7.21 4.68 3.36 2.63 3.56 10.66 9.32 8.23 7.47 8.34
OccWorld-S[zheng2024occworld] Camera None Pred. 0.27 0.28 0.26 0.24 0.26 4.32 5.05 5.01 4.95 5.00
RenderWorld-S[yan2025renderworld] Camera None Pred. - 2.83 2.55 2.37 2.58 - 14.61 13.61 12.98 13.73
COME-S[shi2025come] Camera None Pred. - 25.57 18.35 13.41 19.11 - 45.36 37.06 30.46 37.63
OccWorld-D[zheng2024occworld] Camera Occ Pred. 18.63 11.55 8.10 6.22 8.62 22.88 18.90 16.26 14.43 16.53
OccWorld-F[zheng2024occworld] Camera Occ Pred. 20.09 8.03 6.91 3.54 6.16 35.61 23.62 18.13 15.22 18.99
OccLLaMA-F[wei2024occllama] Camera Occ Pred. 37.38 10.34 8.66 6.98 8.66 38.92 25.81 23.19 19.97 22.99
DFIT-OccWorld-F[zhang2024efficient] Camera Occ Pred. - 13.38 10.16 7.96 10.50 - 19.18 16.85 15.02 17.02
DTT-F[xu2025delta] Camera Occ Pred. 43.52 24.87 18.30 15.63 19.60 54.31 38.98 37.45 31.89 36.11
DOME-F[gu2024dome] Camera Occ GT 75.00 24.12 17.41 13.24 18.25 74.31 35.18 27.90 23.44 28.84
COME-F[shi2025come] Camera Occ GT - 26.56 21.73 18.49 22.26 - 48.08 43.84 40.28 44.07
@@ V-C Control Tasks
PlaNet[hafner2019learning] 5M 469 496 495 945 333/20
Dreamer[hafner2020dream] 5M 935 895 499 962 823/20
Dreaming[okada2021dreaming] 500k 905 566 762 469 610/12
TransDreamer[chen2022transdreamer] 2M - 865 - 933 893/4
DreamerPro[deng2022dreamerpro] 1M 873 897 811 - 857/6
MWM[seo2023masked] 1M - 670 - 690/7
HRSSM[sun2024learning] 500k 910 - 960 - 938/3
DisWM[wang2025disentangled] 1M 960 820 - 920 879/5
@@ Methods
Criteria VidMan[wen2024vidman] ManiGaussian[lu2024manigaussian] ManiGaussian++[yu2025manigaussian++] DreMa[barcellona2025dream] TesserAct[zhen2025tesseract]
@@
UniAD[hu2023planning] Camera Map & Box & Motion & Tracklets & Occ 0.48 0.96 1.65 1.03 0.05 0.17 0.71 0.31
UniAD+DriveWorld[min2024driveworld] Camera Map & Box & Motion & Tracklets & Occ 0.34 0.67 1.07 0.69 0.04 0.12 0.41 0.19
GenAD[zheng2024genad] Camera Map & Box & Motion 0.36 0.83 1.55 0.91 0.06 0.23 1.00 0.43
FSDrive[zeng2025futuresightdrive] Camera Map & Box & QA 0.40 0.89 1.60 0.96 0.07 0.12 1.02 0.40
OccWorld-T[zheng2024occworld] Camera Semantic LiDAR 0.54 1.36 2.66 1.52 0.12 0.40 1.59 0.70
Doe-1[zheng2024doe] Camera QA 0.50 1.18 2.11 1.26 0.04 0.37 1.19 0.53
SSR[li2025navigation] Camera None 0.24 0.65 1.36 0.75 0.00 0.10 0.36 0.15
OccWorld-S[zheng2024occworld] Camera None 0.67 1.69 3.13 1.83 0.19 1.28 4.59 2.02
Epona[zhang2025epona] Camera None 0.61 1.17 1.98 1.25 0.01 0.22 0.85 0.36
RenderWorld[yan2025renderworld] Camera None 0.48 1.30 2.67 1.48 0.14 0.55 2.23 0.97
Drive-OccWorld[yang2025driving] Camera None 0.32 0.75 1.49 0.85 0.05 0.17 0.64 0.29
OccWorld-D[zheng2024occworld] Camera Occ 0.52 1.27 2.41 1.40 0.12 0.40 2.08 0.87
OccWorld-F[zheng2024occworld] Camera Occ 0.45 1.33 2.25 1.34 0.08 0.42 1.71 0.73
OccLLaMA-F[wei2024occllama] Camera Occ 0.38 1.07 2.15 1.20 0.06 0.39 1.65 0.70
DTT-F[xu2025delta] Camera Occ 0.35 1.01 1.89 1.08 0.08 0.33 0.91 0.44
DFIT-OccWorld-V[zhang2024efficient] Camera Occ 0.42 1.14 2.19 1.25 0.09 0.19 1.37 0.55
NeMo[huang2024neural] Camera Occ 0.39 0.74 1.39 0.84 0.00 0.09 0.82 0.30
OccWorld-O[zheng2024occworld] Occ None 0.43 1.08 1.99 1.17 0.07 0.38 1.35 0.60
OccLLaMA-O[wei2024occllama] Occ None 0.37 1.02 2.03 1.14 0.04 0.24 1.20 0.49
RenderWorld-O[yan2025renderworld] Occ None 0.35 0.91 1.84 1.03 0.05 0.40 1.39 0.61
DTT-O[xu2025delta] Occ None 0.32 0.91 1.76 1.00 0.08 0.32 0.51 0.30
DFIT-OccWorld-O[zhang2024efficient] Occ None 0.38 0.96 1.73 1.02 0.07 0.39 0.90 0.45
  • 1 Note: Results are reported following the UniAD [hu2023planning] protocol. Method variants are denoted by their input source: O (ground-truth), V (camera-predicted), D (TPVFormer), F (FBOCC), etc.
@@
VI Challenges and Trends
This section reviews the open challenges and emerging directions for world models in embodied AI. We discuss them across three dimensions: §VI-A Data & Evaluation, §VI-B Computational Efficiency, and §VI-C Modeling Strategies.
VI-A Data & Evaluation
@@
Evaluation practices face similar limitations. Metrics such as FID and FVD emphasize pixel fidelity while ignoring physical consistency, dynamics, and causality. Recent benchmarks, such as EWM-Bench [yue2025ewmbench], introduce new measures but remain task-specific and lack cross-domain standards.
Future Directions.  Recent initiatives such as OpenDV-2K [yang2024generalized] and VideoMix22M [assran2025v] highlight the growing focus on large-scale pretraining and broader modality coverage, yet resources remain fragmented and domain specific. Future work should prioritize constructing unified multimodal, cross-domain datasets to enable transferable pretraining, while advancing evaluation frameworks that move beyond perceptual realism to assess physical consistency, causal reasoning, and long-horizon dynamics.
VI-B Computational Efficiency